OpenStudy (anonymous):

Al Resse and Marci Rey received a total of $210,000 from a deceased relative's estate. They decided to put away $50,000 in a trust for their child and divide the remainder into 1/4 for Marci and 3/4 for Al. How much will Al receive?

OpenStudy (anonymous):

do anyone know if i did this right 210,000 - 50,000 = 160,000 1/4 x 160,000 = $40,000 for Marci 3/4 x 160,000 = $120,000 for Al ]

OpenStudy (anonymous):

yea, i think you did right, cause when you add them together back, 40 000 + 120 000 + 50 000 = 210 000

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

:D